the total value of the pennies and nickels was $14.50. hala counted the coins and found there was 450 in totalin all. how many of each type of coin did she have.
Answer by tutorcecilia(2144)

Use the standard form of a line:
Ax+By=C
Let .01=value of the pennies
Let .05=value of the nickels
.
(.01)(x)+ (.05)(450-x)=14.50
.01x+22.50-.05x=14.50 [Solve for x]
.01x-.05x=-22.50+14.50
-.04x=-8
x=200 pennies
450-200=250 nickels
.
Check by plugging the values back into the original equation:
Ax+By=C
(.01)(200)+(.05)(250)=$14.50
14.50=14.50 [Checks out]